OpenStudy (raden):

How to calculate the sum : 1/sqrt(10000) + 1/sqrt(10001) + 1/sqrt(10002) ... + 1/sqrt(1000000)

OpenStudy (turingtest):

\[\Large\sum_{i=10^4}^{10^6}\frac1{\sqrt i}\]???

OpenStudy (raden):

yea, like that...

OpenStudy (anonymous):

maybe they mean to find floor of this !!

OpenStudy (raden):

but no said to find floor of this question..

OpenStudy (raden):

hmmm... but i want to know using algebraic to solve it give me idea mukushla

OpenStudy (anonymous):

yeah this problem wants the floor... http://www.ocf.berkeley.edu/~wwu/cgi-bin/yabb/YaBB.cgi?board=riddles_medium;action=print;num=1136228390

OpenStudy (anonymous):

integral part means floor

OpenStudy (raden):

ohhh... i just know that "integral part means floor" but i dont look its integral process??

OpenStudy (anonymous):

actually \[{1\over\sqrt{k+1} + \sqrt{k}} < {1\over 2\sqrt{k}} < {1\over\sqrt{k-1} + \sqrt{k}}\]

OpenStudy (anonymous):

rationalize the denum for LHS and RHS of inequality u have\[\sqrt{k+1} - \sqrt{k} < {1\over 2\sqrt{k}} <\sqrt{k} - \sqrt{k-1}\]

OpenStudy (anonymous):

apply the sigma

OpenStudy (raden):

Ok i got it... i think this problem can be solved with integral or simplify one by one that fractions lol... thanks mukushla

OpenStudy (anonymous):

very welcome